OpenStudy (anonymous):

PLEASE HELP! PROBLEM ON RELATED RATES :) A ball is located at point Q on top of a flagpole 10m above the ground. a light is being raised vertically at the rate of 5 m/s casting the shadow of Q on the ground. how fast is the shadow moving along the ground when the light is 50m high?

OpenStudy (anonymous):

where is the light initially?

OpenStudy (anonymous):

20 meters from the flagpole! coming from the top

OpenStudy (amistre64):

dL/dt = 5 ....dh/dt = 0 dx/dt =? dx/dt = dx/dL * dx/dt find dx/dL ?? maybe not tho.....

OpenStudy (anonymous):

is there an initial height of the light?

OpenStudy (amistre64):

ahhh...... the shadow moving problem in all the textbooks :)

OpenStudy (anonymous):

well you need to cross multiply and divide. so it would be 20m/50m by 5m/x.... that would be 50x by 100... divide 50 by 100 to get x alone and you get x=2

OpenStudy (amistre64):

thats a rough drawing by the way :)

OpenStudy (amistre64):

h/(20+x) = 10/(x) if that helps :)

OpenStudy (anonymous):

lol sorry thats what was given to me :p

OpenStudy (amistre64):

y = (200+10x)/x

OpenStudy (amistre64):

dx/dt = x^2(dy/dt)/-200 ?? maybe

OpenStudy (amistre64):

dy/dy = 5

OpenStudy (amistre64):

x=5

OpenStudy (amistre64):

im gonna take a gander and say... 25(5) = 125 125/-200 = ?? -.0625 ?? if I did it right lol

OpenStudy (amistre64):

i think I got the "equation" wrong at the start....

OpenStudy (anonymous):

lol thank you! ..but i dont understand how can you have dy/dy? :s

OpenStudy (anonymous):

lol so the whole thing is wrong??

OpenStudy (amistre64):

:) fat fingers...little tiny keyboard :)

OpenStudy (anonymous):

LMAO! ohhhh i see (:

OpenStudy (amistre64):

it may be right, but I dont trust it yet.... need to find the equation to derive; then the rest is easy.

OpenStudy (anonymous):

ya but thats the part i cant get! wahh :(

OpenStudy (amistre64):

I am thinking it is the relation of congruent triangles; "y" being the height of the lamp: x = length of shadow. 10/x = y/(20+x) is what I come up with; there is no question about the distance of the light to the tip of the shadow...

OpenStudy (amistre64):

10(20+x) = y(x) (200+10x)/x = y 200/x + 10 = y 200/x = y-10 200/(y-10) = x ??? sounds good

OpenStudy (amistre64):

this relate y and x..... now implicit it with respect to time right?

OpenStudy (anonymous):

yea thats what i gotta do

OpenStudy (anonymous):

and thank youuu! for everything so far btw!

OpenStudy (amistre64):

-200(dy/dt) / (y-10)^2 = dx/dt dy/dt = 5 y = 50;

OpenStudy (amistre64):

no prob; I live to math things up lol

OpenStudy (amistre64):

-0.625 is what i get, and I am almost sure its right... but i get that nagging feeling that I am just being stupid :)

OpenStudy (anonymous):

hahaha but where did u get negative two hundred?

OpenStudy (amistre64):

ok.... step it out.... 200 x = ------- (y-10) (y-10)(0) - (200)(dy/dt) dx/dt = ---------------------- (y-10)^2

OpenStudy (anonymous):

lol im on the phone with my friend trying to figure this out and im reading her what u wrote and we both agreed your funny, just thought id let you know :p but yea we dont understand also where you got the 50 from? (: care to explain?

OpenStudy (amistre64):

-200(dy/dt) dx/dt = ------------- (y-10)^2

OpenStudy (amistre64):

when y=50...its right there :)

OpenStudy (amistre64):

dy/dt is given already as the rate of change in "y"... = 5 m/s

OpenStudy (amistre64):

as y rises; x shrinks.... so its (-) direction..

OpenStudy (anonymous):

ohhh okay! thank you so so much honestly, it means a lot (: (:

OpenStudy (amistre64):

just let me know if im right :)

OpenStudy (anonymous):

i will haha but its the first time im on this site so how can i tell you?

OpenStudy (anonymous):

because i have to check with my teacher first

OpenStudy (amistre64):

shout really loud? :) AMISTRE WAS RIGHT!!!.... or if im wrong it would be: AMISTRE WAS WRONG .... ill feel it either way :)

OpenStudy (anonymous):

LMAOOOOOOOOO!

OpenStudy (amistre64):

that moving shadow thing is in all the calc textbooks If seen, and it always makes me nervous

OpenStudy (anonymous):

lol im sure you got it right though, it seemed to make sense. i found it impossible!

OpenStudy (amistre64):

good luck :)

OpenStudy (anonymous):

thank you again!:)